Give the complete solutions of the twoo problems

1.7) Find a+b+c if phi (x) = a(e^x) + b(e^cx) is a solution to the initial value problem: second derivative of y + first derivative of y - 2y = 0; y(0) = 2, first derivative of y(0) = 1. Here a,b, and c are constants.

1.8) Find the positive constant a(>0) so that the parametric equation: x = at, y= -t^2 is a solution to y= x(dy/dx) + (dy/dx)^2.
1
Expert's answer
2020-06-01T19:58:49-0400

1.7) Given "\phi(x) = a e^x + b e^{cx}" is solution of "y''+y' - 2y = 0, y(0)=2,y'(0)=1".

Hence, "y = ae^x+be^{cx} \implies y' = ae^x+bce^{cx} \ and \ y'' = ae^x+bc^2e^{cx}".

So, "(ae^x+bc^2e^{cx}) + (ae^x+bce^{cx}) - 2(ae^x+be^{cx}) = 0"

"\implies b(c^2+c-2) = 0" _____________(1)

Also, "y(0)=2, y'(0)=1"

"\implies a+b=2" __________________(2)

and "a+bc = 1" ___________________(3)

From equation(1) if "b = 0" , then from equation(2) "a = 2" and from equation (3) "a = 1" which is not possible, so "b\neq 0".

Now, so from equation(1), "c^2+c-2 = 0 \implies (c+2)(c-1) = 0"

"\implies either \ c = 1, \ or \ c = -2."

If "c = 1," then from equation (3) "a+b = 1" which contradict the equation (2), so "c \neq 1".

Only "c = -2" is possible, so from equation (3) "a-2b = 1" _____________(4).

So, by equation (2) - equation(3), we get "c = \frac{1}{3}" and hence "a = \frac{5}{3}" .

So, "a+b+c=2+(-2) = 0".


1.8) Given "x = at, y= -t^2" is solution of "y= x(dy\/dx) + (dy\/dx)^2".

Now, "\frac{dx}{dt} = a, \frac{dy}{dt} = -2t \implies \frac{dy}{dx} = \frac{dy/dt}{dx/dt} = -\frac{2t}{a}".

By putting in given differential equation, we get

"-t^2 = at (\frac{-2t}{a}) + (\frac{-2t}{a})^2 \\
\implies -a^2 t ^2= -2 a^2 t^2 + 4 t^2 \implies a^2 = 4 \\ \implies a = 2"